The nation of Potchatoonie produces hockey pucks, cases of root beer, and sandals. The following table provides data on prices and quantities Instructions: Enter your responses as integer values Pucks Root beer Sandals Year Quantity Price Quanti 2011 2014 Price $18 $25 ty Price Quantit 100 125 $5 $9 300 250 $20 $20 100 110 Assume that 2011 is the base year. Find nominal GDP and real GDP for both years Year 2011 2014 Nominal GDP Real GDP
Answer #1

ANSWER:

1) Nominal gdp in 2011 = quantity of pucks * price of pucks + quantity of root beer * price of root beer + quantity of sandals * price of sandals = 100 * 5 + 300 * 20 + 100 * 18 = 500 + 6,000 + 1,800 = 8,300

2) Real gdp in 2011 = quantity of pucks * price of pucks + quantity of root beer * price of root beer + quantity of sandals * price of sandals = 100 * 5 + 300 * 20 + 100 * 18 = 500 + 6,000 + 1,800 = 8,300

3) Nominal gdp in 2014 = quantity of pucks * price of pucks + quantity of root beer * price of root beer + quantity of sandals * price of sandals = 125 * 9 + 250 * 20 + 110 * 25 = 1,125 + 5,000 + 2,750 = 8,875

4) Real gdp in 2014 = quantity of pucks * price of pucks + quantity of root beer * price of root beer + quantity of sandals * price of sandals = 125 * 5 + 250 * 20 + 110 * 18 = 625 + 5,000 + 1,980 = 7,605